A man passes road of 250 meter broad in 75 second. What is the speed of man in km/hour? (A)10 (B) 12 (C) 12.5 (D) 15

Grade:10

1 Answers

Arun
5 years ago
It travels 250 meters in 75 second
Hence
Velocity = 250/75 = 10/3 m/sec
Hence
Velocity in km/hr = 10/3 *18/5 = 180/15 = 12 km/hr